# Zhang, Liu, and Zhao

### v4

@@
Focusing on stochastic programming (SP) with covariate information, this paper proposes an empirical risk minimization (ERM) method embedded within a nonconvex piecewise affine decision rule (PADR), which aims to learn the direct mapping from features to optimal decisions. We establish the nonasymptotic consistency result of our PADR-based ERM model for unconstrained problems, which illustrates the role of piece number in balancing the trade-off between the approximation and estimation errors. To solve the nonconvex and nondifferentiable ERM problem, we develop an enhanced stochastic majorization-minimization (ESMM) algorithm and establish the convergence to (composite strong) directional stationarity along with convergence rate analysis. We show that the proposed PADR-based ERM method applies to a broad class of nonconvex SP problems with theoretical consistency guarantees and computational traceability. Our numerical study demonstrates the effectiveness of PADR-based ERM methods with comparisons to state-of-the-art approaches under various settings, with significantly lower costs, less computation time, and robustness to feature dimensions and nonlinearity of the underlying dependency.
@@ 1 Introduction
With the fast-growing real-world applications and availability of large datasets, the field of stochastic programming (SP) has flourished with many theoretical and algorithmic advances (Birge and Louveaux 2011, Shapiro et al. 2021), in which the empirical probability distribution based on historical scenarios plays a central role in approximating the true distribution. However, it is common in practice that some covariate information (also referred to as contextual information or features) that may affect the random variable could be observed before the decision is made. In such situations, the empirical probability distribution and classical SP methodologies ignoring the covariate information leads to the poor performance. Given the covariate information $X=x\in\mathbb{R}^{p}$, an unconstrained SP is formulated as follows,
@@ 1.1 Problem setting
where $\mathcal{F}$ contains all measurable functions mapping from the feature space $\mathcal{X}$ to the decision space $\mathbb{R}^{d}$. By the interchangeability principle (Shapiro et al. 2021, Theorem 7.80), if a mapping $f^{*}\in\mathcal{F}$ is optimal to problem (2), then the decision $f^{*}(x)$ is optimal to (1) for almost every $x\in\mathcal{X}$. In general, problem (2) is intractable due to the infinite-dimensional function space and the unknown joint distribution of $X$ and $Y$. It is common to consider an approximation method which constructs an empirical risk minimization (ERM) problem based on a dataset $\Xi\coloneqq\{(x^{s},y^{s})\}_{s=1}^{n}$ with the decision rule $f$ restricted to a parametric DR class $\mathcal{H}$ (also known as hypothesis class):
@@
where $\theta\coloneqq\left(\alpha^{k},a^{k},\beta^{k},b^{k}\right)_{k=1}^{K}\in\mathbb{R}^{2K(p+1)}$ and $[K]\coloneqq\{1,\dots,K\}$. The PADR of such a form is beneficial because it requires minimal parameter tuning in multi-dimensional cases without the need to specify breakpoints for piecewise linear parts in each dimension, and it can asymptotically approximate any continuous functions according to Royset (2020). Throughout the theoretical analysis in this paper, we assume that the numbers of pieces in the two max-affine components of (4) are equal to simplify the exposition, which could be easily extended to the case with different numbers of pieces.
@@
In the rest of this section, we give a brief literature review and present the distinguished features of the proposed PADR method over other relative work. We will also summarize major contributions of the present paper.
@@ 1.2 Literature review
However, as claimed by Bertsimas and Koduri (2021), the prescriptive methods are highly affected by the data with features close to the current observation and thus suffers from the curse of dimensionality. The DR-based ERM framework presented in (3), which is another one-step approach in the integrated scheme, is relieved from the aforementioned issue due to its functional approximation in nature. For constrained SP with general nonlinear constraints, the current study of decision rule methods could only guarantee feasibility with high probability, and the predict-then-optimize framework may be more suitable for solving the contextual problems when the feasibility is essential to decision-makers.
Several works on DR-based ERM methods are most relevant to this paper. Ban and Rudin (2018) employ LDR for the Newsvendor problem with nonasymptotic optimality guarantees under the assumption of the linear demand model. Notz and Pibernik (2021) and Bertsimas and Koduri (2021) study decision rules in reproducing kernel Hilbert spaces (RKHS), which we refer to as RKHS-DR, and provide asymptotic optimality guarantees for convex SP problems. Compared to the above existing DR methods, given the asymptotic approximation of PA functions to any continuous functions, the PADR method is applicable to more general decision-making problems with nonconvex cost function or nonlinear dependency with asymptotic optimality.
@@
Besides the classical LDR, several generalized LDRs are proposed for standard SP problems which lift random vectors into a higher-dimensional probability space to describe nonlinear relations (Chen et al. 2007, Bampou and Kuhn 2011, Georghiou et al. 2015). However, they tend to have unsatisfactory performance without cautious prior designs especially when the random vector is high-dimensional; moreover, they usually lack theoretical guarantees on approximation accuracy except for the posterior estimation provided in Kuhn et al. (2011).
@@
Regarding solving the PADR-based ERM problem with potentially large sample size $n$, it is prohibitively challenging to obtain a global minimizer to such a coupled nonconvex and nondifferentiable optimization problem. As provided by Bertsimas and Georghiou (2015, Section 2.4), the PADR-based ERM problem can be formulated as a mixed-integer nonlinear program using Big-M method with additional $2ndK$ binary variables to describe the PADR of interest. For a PADR-based ERM problem of the modest size with data size $n=1000$, the obtained mixed-integer nonlinear program could have $10^{4}\sim 10^{5}$ binary variables and the state-of-the-art solvers are hard to solve within a few hours to achieve an approximately optimal solution.
@@
Instead of seeking global optimality, majorization-minimization (MM) methods provide a unified framework that is practically efficient and theoretically convergent for solving the PADR-based ERM problem with the finite-sum composite objective. MM methods iteratively construct convex subproblems with an upper surrogate objective function for obtaining “stationary” solutions. According to variational analysis by Rockafellar and Wets (2009), there are many kinds of stationary concepts which differ in the ability to characterize necessary optimality conditions for local/global optimality. As pointed in Cui and Pang (2021, Chapter 7), MM algorithms may converge to different types of stationary solutions depending on the properties of surrogate and objective functions. Algorithms with convergence to a stronger stationarity is favorable not only theoretically but also computationally because the obtained solution from numerical implementation with random initialization has the greater potential to be a global optimum.
@@
Difference-of-convex (DC) algorithms, which is a special MM method, have been extensively studied over the last several decades (Le Thi and Pham Dinh 2018) and can compute a d(irectional)-stationary point for DC programs with a differentiable concave component. When the cost function $\varphi(\cdot,y)$ is convex or piecewise affine, the PADR-based ERM problem can be formulated as a nondifferentiable DC program for which DC algorithms including Tao and An (1997) and Lipp and Boyd (2016) are applicable. However, due to the nondifferentiable concave component, DC algorithms for the ERM problem converge to a critical point whose defining condition is dependent on the dc decomposition form and is weaker than d(irectional)-stationarity. For nondifferentiable DC or composite DC problems, Pang et al. (2016) and Cui et al. (2018) propose an enhanced surrogation technique that iteratively solves convex subproblems that are constructed based on $\varepsilon$-active pieces of the concave component. However, the enhanced MM algorithm for large-scale ERM problems requires iteratively solving finite-sum convex subproblems over all the data, which raises the fundamental issue of computational efficiency. Sampling-based algorithms for stochastic DC programs have been recently studied in Nitanda and Suzuki (2017), Xu et al. (2019), Le Thi et al. (2022) addressing smooth or partially smooth objectives, and in Liu et al. (2022) addressing nonconvex and nonsmooth compound SP to compute critical solutions. There still lacks a comprehensive study on whether the $\varepsilon$-active set technique is amenable for sampling-based methods to enhance the convergence properties for nondifferentiable stochastic DC programs and PADR-based ERM problems.
@@ 1.3 Contributions
We summarize our major contributions as follows.
1. We propose a PADR-based ERM method for a general class of SP problems with covariate information, where the cost function could be nonconvex and nondifferentiable. We quantify the approximation accuracy of the PA function class to any Lipschitz continuous function, which to the best of our knowledge is the first quantification result for such a general case. Without any parametric assumption on the optimal decision rule or the latent dependency, we establish the nonasymptotic consistency guarantee with the excess risk bound of our proposed framework, which illustrates the role of piece number in balancing the approximation and estimation errors.
2. For efficiently solving the PADR-based ERM problem, we develop an enhanced sampling-based MM (ESMM) algorithm with the natural combination of sequential sampling and the enhanced technique over the $\varepsilon$-active set. With sequential sampling, the ESMM algorithm effectively reduces computational costs by solving subproblems of small sizes in the early iterations. We show that under the proper control of increasing sample rate, ESMM achieves the asymptotic convergence in terms of composite strong d-stationarity, which is strictly stronger than d-stationarity and thus has greater potential to be a global minimizer. Numerical results show that the ESMM algorithm could speed up to 4 times faster than the basic MM method.
We also would like to highlight several distinguishing features of the proposed PADR method for contextual SP: (a) from the modeling aspect, the PADR method is suitable for contextual SP problems with nonlinear or nonconvex cost functions. All other relevant works on contextual SP with asymptotic optimality are restricted to linear programs or convex programs. For instance, the RKHS-based DR method by Bertsimas and Koduri (2021) requires convexity of the cost function for asymptotic optimality. The PADR method opens up the application to nonconvex contextual SP problems; (b) from the theoretical aspect, PADR provides a compact structure to asymptotically approximate non-parametric decision rule with minimal hyperparameter tuning. Though PA hypothesis class is not new in classical SP, we provide the first quantification of the approximation error and the excess risk bound of the data-driven PADR model for general nonconvex SP problems. Except for the LDR method, all other methods for contextual SP do not provide the excess risk bound, including the RKHS-based DR in Bertsimas and Koduri (2021) and Notz and Pibernik (2021); (c) regarding to the computational aspect, the proposed ESMM algorithm is developed for efficiently solving the PADR-based ERM problem of the large scale. Under the suitable sampling scheme, we show that the ESMM method significantly improves computational efficiency and enhances the convergence property compared to the existing MM methods in the literature; (d) moreover, the PADR method possesses several prominent numerical advantages. Numerous experiments indicate that it requires significantly fewer data points for satisfactory performance, thus is particularly suitable for cases with insufficient data sets. Additionally, it yields superior performance in terms of the test costs with both synthetic and real data sets when the feature dimension is less than 20. The nonlinearity level of the demand model has almost no effect on the performance of the PADR method, while other methods suffer significantly from the high level of nonlinearity.
@@ 2.1 Approximation error of piecewise affine function class
The approximation errors of the PA function class to convex and DC functions are quantified in Balazs et al. (2015) and Siahkamari et al. (2020) respectively. PA functions which are of “nonparametric” nature are known to asymptotically approximate Lipschitz continuous functions (Royset 2020); nevertheless, the approximation error of PA functions with finite pieces remains unsolved. For the class of Lipschitz continuous functions that allow a more general setting for SP problems of interest, we provide the approximation error bound of the PA function class, which to the best of our knowledge is the first quantification result for such a general case.
@@
[…]
@@ 2.2 Nonasymptotic consistency of PADR-based ERM
The boundedness of $\mathcal{X}$ in Assumption (A2) is common in practical applications. Under Assumption (A2), the optimal DR $f^{*}$ in Assumption (A4) is bounded on $\mathcal{X}$, and thus $f^{*}\in\mathcal{F}_{L_{z},M_{z}}$ for some $L_{z}>0$ and $M_{z}>0$.
@@
[…]
@@
Under Assumptions (A1)-(A3), with $M_{\mu}\coloneqq C_{\varphi}+L_{\varphi}\left(2\sqrt{d}\mu(p\bar{X}+1)+\left\lVert z_{0}\right\rVert\right)$, the following holds with probability at least $1-\delta$ over the draw of $\Xi$,
$\max_{f\in{\mathcal{H}}_{\mu}^{K}}\left\lvert R(f)-\widehat{R}_{\Xi}(f)\right\rvert\leq\frac{8\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+2M_{\mu}\sqrt{\frac{2dK(p+1)\log n}{n}}+\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$
@@
Under Assumptions (A1)-(A4), with $\mu\geq\max\left\{\frac{1}{2}L_{z}K^{1/p},\frac{1}{4}pL_{z}\bar{X}K^{1/p}+\frac{1}{2}M_{z}\right\}$ and $M_{\mu}\coloneqq C_{\varphi}+L_{\varphi}\left(2\sqrt{d}\mu(p\bar{X}+1)+\left\lVert z_{0}\right\rVert\right)$, with probability at least $1-\delta$ over the draw of $\Xi$, we have
@@
$+\frac{16\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+4M_{\mu}\sqrt{\frac{2dK(p+1)\log n}{n}}+2\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$
@@
where $\widetilde{\mathcal{O}}(\cdot)$ suppresses logarithmic factors. It can be seen that the sample size $n$ only affects the estimation error, whereas $K$, the number of max-affine pieces for PA functions in ${\mathcal{H}}_{\mu}^{K}$, plays the central role of balancing the two error terms. As $K$ grows, the PA function class becomes larger, so the approximation error decreases due to the stronger approximation ability of ${\mathcal{H}}_{\mu}^{K}$, and the estimation error increases due to the more pronounced overfitting issue. With an appropriate choice of $K$ that balances these two errors, we can derive the optimal excess risk bound. Furthermore, with the existence of a convex or DC optimal DR, this bound can be improved based on the approximation bounds $\mathcal{O}(K^{-2/p})$ of ${\mathcal{H}}_{\mu}^{K}$ from Balazs et al. (2015) and Siahkamari et al. (2020). The extensions are summarized in the following corollary.
@@
The excess risk bound (8) illustrates the role of $K$ in balancing the two errors, which is further evidenced by numerical experiments summarized in the table below. Table 2.2 reports the validation and training costs provided by the PADR-based ERM method in solving a Newsvendor problem with the sine plus max-affine demand model (19) (details are given in Section 4.1), under multiple choices of $K_{1}$ and $K_{2}$ which represent the number of pieces for the two max-affine components in the PADR respectively. The validation cost, which is the empirical cost of the obtained PADR solution over the validation data set of size 1000, can be regarded as the sample average approximation of $R(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})$. Noticing that $\inf_{f\in\mathcal{F}}R(f)$ is irrelevant to the PA hypothesis class, the performance of excess risk could be revealed through the pattern of validation costs in the table below. For $K_{1}$ and $K_{2}$ being small, the validation cost decreases along with increasing values of $K_{1}$ and $K_{2}$ which may be credited to the decreasing approximation error. For $K_{1}$ and $K_{2}$ being larger than 4, the validation cost slightly increases, suggesting that as the PADR model becomes more complex, the increasing estimation error gradually dominates the decreasing approximation error. Hence, the best validation costs shall be achieved with moderate values of $K_{1}$ and $K_{2}$, which are usually considerably smaller than the ones implied by the theoretical results.
@@
Validation and training costs for basic Newsvendor with demand model (19) $K_{1}\backslash K_{2}$ 1 2 3 4 5 6 1 24.06 (24.09) 23.62 (23.57) 23.60 (23.63) 23.64 (23.55) 23.72 (23.66) 23.64 (23.62) 2 8.68 (8.73) 7.01 (7.09) 7.01 (7.09) 7.02 (7.08) 7.00 (7.10) 7.02 (7.09) 3 7.82 (7.95) 5.41 (5.53) 5.45 (5.51) 5.44 (5.51) 5.43 (5.49) 5.42 (5.49) 4 7.29 (7.47) 4.26 (4.27) 4.25 (4.22) 4.28 (4.29) 4.29 (4.21) 4.22 (4.23) 5 7.26 (7.46) 4.26 (4.27) 4.25 (4.24) 4.23 (4.21) 4.25 (4.21) 4.27 (4.22) 6 7.28 (7.47) 4.27 (4.24) 4.17 (4.23) 4.16 (4.23) 4.24 (4.27) 4.22 (4.20) The PADR is obtained by Algorithm 1 for solving the PADR-based ERM model. Numbers listed in brackets represent training costs, and the lowest costs are highlighted in bold and red text.

### v5

@@
Focusing on stochastic programming (SP) with covariate information, this paper proposes an empirical risk minimization (ERM) method embedded within a nonconvex piecewise affine decision rule (PADR), which aims to learn the direct mapping from features to optimal decisions. We establish the nonasymptotic consistency result of our PADR-based ERM model for unconstrained problems, which illustrates the role of piece number in balancing the trade-off between the approximation and estimation errors. To solve the nonconvex and nondifferentiable ERM problem, we develop an enhanced stochastic majorization-minimization (ESMM) algorithm and establish the convergence to (composite strong) directional stationarity along with convergence rate analysis. We show that the proposed PADR-based ERM method applies to a broad class of nonconvex SP problems with theoretical consistency guarantees and computational tractability. Our numerical study demonstrates the effectiveness of PADR-based ERM methods with comparisons to state-of-the-art approaches under various settings, with significantly lower costs, less computation time, and robustness to feature dimensions and nonlinearity of the underlying dependency.
@@ 1 Introduction
With the fast-growing real-world applications and availability of large datasets, the field of stochastic programming (SP) has flourished with many theoretical and algorithmic advances (Birge and Louveaux 2011, Shapiro et al. 2021a), in which the empirical probability distribution based on historical scenarios plays a central role in approximating the true distribution. However, it is common in practice that some covariate information (also referred to as contextual information or features) that may affect the random variable could be observed before the decision is made. In such situations, the empirical probability distribution and classical SP methodologies ignoring the covariate information lead to the poor performance. Given the covariate information $X=x\in\mathbb{R}^{p}$, an unconstrained SP is formulated as follows,
@@ 1.1 Problem setting
where $\mathcal{F}$ contains all measurable functions mapping from the feature space $\mathcal{X}$ to the decision space $\mathbb{R}^{d}$. By the interchangeability principle (Shapiro et al. 2021a, Theorem 7.80), if a mapping $f^{*}\in\mathcal{F}$ is optimal to problem (2), then the decision $f^{*}(x)$ is optimal to (1) for almost every $x\in\mathcal{X}$. In general, problem (2) is intractable due to the infinite-dimensional function space and the unknown joint distribution of $X$ and $Y$. It is common to consider an approximation method which constructs an empirical risk minimization (ERM) problem based on a dataset $\Xi\coloneqq\{(x^{s},y^{s})\}_{s=1}^{n}$ with the decision rule $f$ restricted to a parametric DR class $\mathcal{H}$ (also known as hypothesis class):
@@
where $\theta\coloneqq\left(\alpha^{k},a^{k},\beta^{k},b^{k}\right)_{k=1}^{K}\in\mathbb{R}^{2K(p+1)}$ and $K]\coloneqq\{1,\dots,K\}$. The PADR of such a form is beneficial because it requires minimal parameter tuning in multi-dimensional cases without the need to specify breakpoints for piecewise linear parts in each dimension, and it can asymptotically approximate any continuous function according to Royset (2020). Throughout the theoretical analysis in this paper, we assume that the numbers of pieces in the two max-affine components of (4) are equal to simplify the exposition, which could be easily extended to the case with different numbers of pieces.
@@
In the rest of this section, we give a brief literature review and present the distinguished features of the proposed PADR method over other related work. We will also summarize the major contributions of the present paper.
@@ 1.2 Literature review
However, as claimed by Bertsimas and Koduri (2021), the prescriptive methods are highly affected by the data with features close to the current observation and thus suffer from the curse of dimensionality. The DR-based ERM framework presented in (3), which is another one-step approach in the integrated scheme, is relieved from the aforementioned issue due to its functional approximation in nature. For constrained SP with general nonlinear constraints, the current study of decision rule methods could only guarantee feasibility with high probability, and the predict-then-optimize framework may be more suitable for solving the contextual problems when feasibility is essential to decision-makers. We refer interested readers to Qi and Shen (2022) and Sadana et al. (2025) for comprehensive reviews on recent advances in both two-stage and integrated approaches in contextual SP.
@@
Several works on DR-based ERM methods are most relevant to this paper. Ban and Rudin (2018) employ LDR for the Newsvendor problem with nonasymptotic optimality guarantees under the assumption of the linear demand model. Besides the classical LDR, several generalized LDRs are proposed for standard SP problems which lift random vectors into a higher-dimensional probability space to describe nonlinear relations (Chen et al. 2007, Bampou and Kuhn 2011, Georghiou et al. 2015). However, they tend to have unsatisfactory performance without cautious prior designs especially when the random vector is high-dimensional; moreover, they usually lack theoretical guarantees on approximation accuracy except for the posterior estimation in Kuhn et al. (2011).
Notz and Pibernik (2021) and Bertsimas and Koduri (2021) study decision rules in reproducing kernel Hilbert spaces (RKHS), which we refer to as RKHS-DR, and provide asymptotic optimality guarantees for convex SP problems. With the rise of deep neural networks (NN), several studies apply NN-based decision rules (NN-DR) in practical applications, such as inventory management (Zhang et al. 2020, Oroojlooyjadid et al. 2020, Qi et al. 2023) and portfolio optimization (Zhang et al. 2020). NN-DRs are powerful in approximating the underlying relationship between decisions and covariate information; however, they may face numerical challenges for complex decision-making problems and also lack theoretical consistency guarantees. Compared to the above existing DR methods, given the asymptotic approximation of PA functions to any continuous function, we show that the PADR method is applicable to a class of more general decision-making problems with nonconvex cost function with both numerical efficiency and theoretical consistency.
@@
Regarding solving the PADR-based ERM problem with potentially large sample size $n$, it is prohibitively challenging to obtain a global minimizer to such a coupled nonconvex and nondifferentiable optimization problem. As provided by Bertsimas and Georghiou (2015, Section 2.4), the PADR-based ERM problem can be formulated as a mixed-integer nonlinear program using Big-M method with additional $2ndK$ binary variables to describe the PADR of interest. For a PADR-based ERM problem of the modest size with data size $n=1000$, the obtained mixed-integer nonlinear program could have $10^{4}\sim 10^{5}$ binary variables and state-of-the-art solvers find it hard to solve within a few hours to achieve an approximately optimal solution.
@@
Instead of seeking global optimality, majorization-minimization (MM) methods provide a unified framework that is practically efficient and theoretically convergent for solving the PADR-based ERM problem with the finite-sum composite objective. MM methods iteratively construct convex subproblems with an upper surrogate objective function for obtaining “stationary” solutions. According to variational analysis by Rockafellar and Wets (2009), there are many kinds of stationary concepts which differ in the ability to characterize necessary optimality conditions for local/global optimality. As pointed out in Cui and Pang (2021a, Chapter 7), MM algorithms may converge to different types of stationary solutions depending on the properties of surrogate and objective functions. Algorithms with convergence to a stronger stationarity are favorable not only theoretically but also computationally because the obtained solution from numerical implementation with random initialization has the greater potential to be a global optimum.
@@
Difference-of-convex (DC) algorithms, which is a special MM method, have been extensively studied over the last several decades (Le Thi and Pham Dinh 2018) and can compute a d(irectional)-stationary point for DC programs with a differentiable concave component. When the cost function $\varphi(\cdot,y)$ is convex or piecewise affine, the PADR-based ERM problem can be formulated as a nondifferentiable DC program for which DC algorithms including Tao and An (1997) and Lipp and Boyd (2016) are applicable. However, due to the nondifferentiable concave component, DC algorithms for the ERM problem converge to a critical point whose defining condition is dependent on the DC decomposition form and is weaker than d(irectional)-stationarity. For nondifferentiable DC or composite DC problems, Pang et al. (2016) and Cui et al. (2018a) propose an enhanced surrogation technique that iteratively solves convex subproblems that are constructed based on $\varepsilon$-active pieces of the concave component. However, the enhanced MM algorithm for large-scale ERM problems requires iteratively solving finite-sum convex subproblems over all the data, which raises the fundamental issue of computational efficiency. Sampling-based algorithms for stochastic DC programs have been recently studied in Nitanda and Suzuki (2017), Xu et al. (2019), Le Thi et al. (2022) addressing smooth or partially smooth objectives, and in Liu et al. (2022a) addressing nonconvex and nonsmooth compound SP to compute critical solutions. There is still a lack of a comprehensive study on whether the $\varepsilon$-active set technique is amenable for sampling-based methods to enhance the convergence properties for nondifferentiable stochastic DC programs and PADR-based ERM problems.
@@ 1.3 Contributions
Our work contributes to the literature on contextual SP along three major dimensions:
1. Modeling aspect. We propose a PADR-based ERM method for a general class of SP problems with covariate information, where the cost function could be nonconvex and nondifferentiable. PADR provides a compact structure to approximate non-parametric decision rules. Though PA hypothesis class is not new in classical SP, we provide the first quantification of the approximation error and the excess risk bound of the data-driven PADR model for general nonconvex contextual SP problems. The excess risk bound illustrates the role of the piece numbers $K_{1}$ and $K_{2}$ in balancing the approximation and estimation errors, which is further evidenced by numerical experiments. In contrast, previous DR methods for contextual SP provide asymptotic guarantees only limited to convex or linear settings. The PADR method thereby provides a rigorous framework to solve nonconvex contextual SP problems with theoretical guarantees.
2. Computational aspect. By exploiting the compact max-affine structure of PADR, we develop the ESMM algorithm to solve the highly nonconvex and nonsmooth PADR-based ERM problem, with the natural combination of sequential sampling and the enhanced technique over the $\varepsilon$-active set within the majorization-minimization framework. We analyze the asymptotic convergence of ESMM in terms of composite strong d-stationarity which is strictly stronger than d-stationarity, the sharpest type of first-order stationary solutions for general nonconvex optimization. Furthermore, with the design of a suitable sampling scheme, the ESMM method significantly improves computational efficiency compared to existing MM approaches. Numerical results show that the ESMM algorithm for PADR scales well to large problem sizes (e.g., up to 80 pieces in PADR) and could be up to 4 times faster than the basic MM method.
3. Empirical aspect. Numerical experiments with comparisons to a variety of prominent methods including NN-DR and RKHS-DR (see Table 4) show that PADR requires fewer data points to achieve satisfactory performance, and thus it is suitable for cases with insufficient data sets. Additionally, it yields superior performance in terms of the test costs with both synthetic and real data sets in various instances including the Newsvendor problem and the product placement problem with feature dimension less than 20. Finally, PADR is more robust to the level of nonlinearity in demand models and consistently obtains satisfactory performance with highly nonlinear demand models, compared to all other benchmarks.
@@ 2.1 Approximation error of piecewise affine function class
The approximation errors of the PA function class to convex and DC functions are quantified in Balazs et al. (2015) and Siahkamari et al. (2020) respectively. PA functions that are of “nonparametric” nature are known to asymptotically approximate Lipschitz continuous functions (Royset 2020); nevertheless, the approximation error of PA functions with finite pieces remains unsolved. For the class of Lipschitz continuous functions that allow a more general setting for SP problems of interest, we provide the approximation error bound of the PA function class, which to the best of our knowledge is the first quantification result for such a general case.
@@
$\mathcal{F}_{L,M}\coloneqq\left\{f:\mathcal{X}\rightarrow\mathbb{R}^{d}\mid\left\lVert f\right\rVert_{\infty}\leq M,\,\text{$f$ is $L$-Lipschitz continuous}\right\},$
@@ 2.2 Nonasymptotic consistency of PADR-based ERM
The boundedness of $\mathcal{X}$ in Assumption (A2) is common in practical applications. Assumption (A4) may appear to be strict, however, in Online Appendix 10, we show that it applies to a broad class of problems. We also provide examples of problems that violate this assumption to further clarify its scope of applicability. Under Assumption (A2), the optimal DR $f^{*}$ in Assumption (A4) is bounded on $\mathcal{X}$, and thus $f^{*}\in\mathcal{F}_{L_{z},M_{z}}$ for some $L_{z}>0$ and $M_{z}>0$.
@@
The approximation error, measuring the gap between the minimal expected costs over the PADR class and the class of all measurable functions, can thus be bounded based on the approximation error bound of PADR class provided in Proposition 2.1. We provide the uniform generalization bound based on the Rademacher complexity theory in Proposition 2.2 which is used to bound the estimation error. Combining the approximation and estimation error bounds, we thus obtain the excess risk bound of PADR-based ERM in Theorem 2.3. The proof of the following two results are provided in Online Appendix 6.2. Note that in the present paper the notation $\left\lVert\cdot\right\rVert$ without a subscript refers to the Euclidean norm.
@@
Under Assumptions (A1)-(A3), with $M_{p,d}(\mu)\coloneqq C_{\varphi}+L_{\varphi}\left(2\sqrt{d}\mu(p\bar{X}+1)+\left\lVert z_{0}\right\rVert\right)$, the following holds with probability at least $1-\delta$ over the draw of $\Xi$,
$\max_{f\in{\mathcal{H}}_{\mu}^{K}}\left\lvert R(f)-\widehat{R}_{\Xi}(f)\right\rvert\leq$ $\ \frac{8\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+2M_{p,d}(\mu)\sqrt{\frac{2dK(p+1)\log n}{n}}$
$\ +\sqrt{2}M_{p,d}(\mu)\sqrt{\frac{\log(2/\delta)}{n}}.$
@@
Under Assumptions (A1)-(A4), with $\mu\geq\max\left\{\frac{1}{2}L_{z}K^{1/p},\frac{1}{4}pL_{z}\bar{X}K^{1/p}+\frac{1}{2}M_{z}\right\}$ and $M_{p,d}(\mu)\coloneqq C_{\varphi}+L_{\varphi}\left(2\sqrt{d}\mu(p\bar{X}+1)+\left\lVert z_{0}\right\rVert\right)$, with probability at least $1-\delta$ over the draw of $\Xi$, we have
@@
$+\frac{16\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+4M_{p,d}(\mu)\sqrt{\frac{2dK(p+1)\log n}{n}}+2\sqrt{2}M_{p,d}(\mu)\sqrt{\frac{\log(2/\delta)}{n}}.$
@@
[…]
@@
The excess risk bound (8) illustrates the role of $K$ in balancing the two errors, which is further evidenced by numerical experiments summarized in the table below. Table 2.2 reports the validation and training costs provided by the PADR-based ERM method in solving a Newsvendor problem with the sine plus max-affine demand model (19) (details are given in Section 4.1), under multiple choices of $K_{1}$ and $K_{2}$ which represent the number of pieces for the two max-affine components in the PADR respectively. The validation cost, which is the empirical cost of the obtained PADR solution over the validation data set of size 1000, can be regarded as the sample average approximation of $R(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})$. Noticing that $\inf_{f\in\mathcal{F}}R(f)$ is irrelevant to the PA hypothesis class, the performance of excess risk could be revealed through the pattern of validation costs in the table below. The training cost decreases consistently as $K_{1}$ and $K_{2}$ increase, indicating the stronger fitting capacity of PADR on the training data. For small values of $K_{1}$ and $K_{2}$, the validation cost also decreases significantly, which can be attributed to the reduction in approximation error. However, this decreasing trend does not hold as $K_{1}$ and $K_{2}$ become large; instead, the validation cost begins to plateau or even slightly increase. For instance, when $K_{2}$ is fixed at 40, the validation cost first decreases to 4.13 (at $K_{1}=5$) and then increases to 4.24 (at $K_{1}=20$). This suggests that as the PADR model becomes more complex, the increasing estimation error gradually dominates the decreasing approximation error. Hence, the best validation costs shall be achieved with moderate values of $K_{1}$ and $K_{2}$, which are usually considerably smaller than the ones implied by the theoretical results.
@@
Validation and training costs for basic Newsvendor with demand model (19) K1 $\backslash$ K2 1 2 3 4 5 10 20 40 1 24.06 (24.09) 23.62 (23.57) 23.61 (23.61) 23.64 (23.55) 23.56 (23.62) 23.55 (23.64) 23.54 (23.62) 23.56 (23.64) 2 8.68 (8.73) 7.01 (7.09) 7.07 (7.00) 7.02 (7.08) 7.08 (6.99) 7.05 (7.02) 7.05 (7.00) 7.05 (7.01) 3 7.95 (7.79) 5.60 (5.40) 5.47 (5.39) 5.40 (5.35) 5.42 (5.34) 5.36 (5.33) 5.41 (5.37) 5.39 (5.35) 4 7.29 (7.47) 4.26 (4.27) 4.25 (4.14) 4.28 (4.29) 4.14 (4.17) 4.13 (4.25) 4.15 (4.15) 4.14 (4.21) 5 7.41 (7.23) 4.21 (4.17) 4.34 (4.14) 4.32 (4.15) 4.17 (4.13) 4.34 (4.16) 4.14 (4.15) 4.11 (4.13) 10 7.41 (7.25) 4.15 (4.18) 4.15 (4.11) 4.10 (4.13) 4.14 (4.14) 4.06 (4.24) 4.04 (4.25) 4.07 (4.14) 20 7.41 (7.24) 4.16 (4.21) 4.15 (4.14) 4.10 (4.14) 4.32 (4.11) 4.04 (4.14) 4.00 (4.15) 4.03 (4.24) 40 7.40 (7.23) 4.15 (4.14) 4.12 (4.15) 4.08 (4.15) 4.11 (4.13) 4.01 (4.24) 3.99 (4.26) 4.02 (4.15) The PADR is obtained by Algorithm 1 for solving the PADR-based ERM model. Numbers listed in brackets represent training costs, and the lowest costs are highlighted in bold and blue text.